) If the marginal propensity to consume is 0.5, find by how much government

Spending would have to rise in order to raise output by Kes 1,000 billion


1
Expert's answer
2021-01-22T12:57:28-0500

"\bold {Answer}"

Government spending should have to rise by "KES \space 500 \space billion"


"\bold {Solution}"

It is given that:

"MPC = 0.5"

"∆Y = KES \space 1,000 \space billion"


Then,

The expenditure multiplier,

"k = \dfrac {1}{1-MPC}"


"= \dfrac {1}{1-0.5}"


"= \dfrac {1}{0.5}"


"= 2"


"\therefore \space ∆G × k = ∆Y"


"=> ∆G × 2 = KES \space 1,000 \space bn"


"=> ∆G = \dfrac {KES \space 1,000 \space bn}{2}"


"= \bold {KES \space 500 \space billion}"


Thus, government spending should rise by KES 500 billion.